OpenStudy (s117):

What is the value of the expression 14 + 6 • 2 – 10 ÷ 5? A. 6 B. 24 C. 0 D. 38

OpenStudy (anonymous):

We use the order of PEMDAS: P: Parenthesis E: Exponent M/D: Multiplication and Division starting from left to right. A/S: Addition and Subtraction starting from left to right. So what is our first step here? @S117

OpenStudy (anonymous):

PEDMAS ^

OpenStudy (s117):

24?

OpenStudy (igreen):

@iambatman Actually it's PEMDAS.. Please Excuse My Dear Aunt Sally But it doesn't really matter..

OpenStudy (igreen):

And yes, 24 is correct.

OpenStudy (anonymous):

Oh jeeze, there's even a saying for it?

OpenStudy (briellshaw):

the answer is 6 want me to show work

OpenStudy (igreen):

Yes, it's very stupid, but whatever..

OpenStudy (anonymous):

Ikr? @iambatman

OpenStudy (igreen):

@BriellShaw Yes please..because 6 is incorrect.

OpenStudy (anonymous):

I'm going to my batcave, idk what's going on with OS today, but I'm done.

OpenStudy (s117):

lel

OpenStudy (briellshaw):

14 + 6 • 2 – 10 ÷ 5? 14+6=20*2=40-10=30/5=6

OpenStudy (briellshaw):

hope it helps

OpenStudy (s117):

it doesnt

OpenStudy (briellshaw):

how not

OpenStudy (briellshaw):

I mean im only a 7th grader it I cant really help much if your above 7th grade

OpenStudy (s117):

because you didnt use pemdas

OpenStudy (s117):

i am to

OpenStudy (briellshaw):

Please Excuse My Dear Aunt Sally

OpenStudy (s117):

yup

OpenStudy (s117):

P: Parenthesis E: Exponent M/D: Multiplication and Division starting from left to right. A/S: Addition and Subtraction starting from left to right.

OpenStudy (briellshaw):

Parentheses exponents multiplication division addition subtraction

OpenStudy (briellshaw):

pemdas

OpenStudy (s117):

@iGreen

OpenStudy (briellshaw):

can you say my answer was the best response

OpenStudy (s117):

no